# Learning out-of-time-ordered correlators with classical kernel methods

### v1

@@ C.2 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{1}$
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear RMSE 6.6797 7.2117 7.2580 7.3482 7.2775 7.2564 7.1498 7.0573 MAE 5.8351 6.4148 6.4226 6.5026 6.3839 6.3326 6.1876 6.0515 $R^{2}$2.6121 -3.7405 -3.5677 -3.5157 -3.3017 -3.1259 -2.8970 -2.6603 Polynomial RMSE 0.9266 1.1954 1.2084 1.4719 1.4644 1.5292 1.5170 1.5602 MAE 0.6546 0.8219 0.8558 1.0416 1.0461 1.0767 1.0694 1.1116 $R^{2}$ 0.9305 0.8697 0.8734 0.8188 0.8258 0.8168 0.8246 0.8211 RBF RMSE 0.7402 0.8833 0.8885 1.1049 1.1179 1.0755 1.0196 0.9878 MAE 0.5409 0.4966 0.4385 0.6916 0.6624 0.6367 0.5657 0.5190 $R^{2}$ 0.9556 0.9289 0.9315 0.8979 0.8985 0.9094 0.9208 0.9283 Laplacian RMSE 0.6841 0.6176 0.6979 0.7291 0.7592 0.8534 0.8078 0.7417 MAE 0.4643 0.3579 0.3771 0.3990 0.4168 0.5653 0.5250 0.4746 $R^{2}$ 0.9621 0.9652 0.9578 0.9555 0.9532 0.9429 0.9503 0.9596 Sigmoid RMSE 1.7673 1.1875 1.9128 1.4774 1.4508 1.3716 2.0234 2.2092 MAE 1.3062 0.7939 1.4378 1.0354 1.0400 0.9645 1.5305 1.7456 $R^{2}$ 0.7472 0.8715 0.6827 0.8175 0.8290 0.8526 0.6879 0.6413 Cosine RMSE 6.7042 7.2079 7.2421 7.3244 7.2756 7.2927 7.1820 7.0846 MAE 5.8541 6.4031 6.3958 6.4589 6.3751 6.3714 6.2191 6.0739 $R^{2}$2.6387 -3.7354 -3.5476 -3.4864 -3.2995 -3.1673 -2.9321 -2.6887 Training Linear RMSE 6.4890 7.0385 7.0937 7.1710 7.1516 7.1017 6.9922 6.8715 MAE 5.5720 6.1532 6.1925 6.2540 6.1792 6.1015 5.9658 5.8112 $R^{2}$2.1828 -3.1641 -3.0659 -2.9840 -2.8636 -2.6929 -2.5129 -2.3154 Polynomial RMSE 0.7307 0.8701 0.9015 1.2510 1.2271 1.2795 1.2360 1.2827 MAE 0.5112 0.6282 0.6547 0.8919 0.8785 0.8960 0.8745 0.9044 $R^{2}$ 0.9596 0.9364 0.9343 0.8788 0.8862 0.8801 0.8902 0.8845 RBF RMSE 0.5713 0.4603 0.1088 0.7159 0.5706 0.6236 0.4932 0.4137 MAE 0.4250 0.3093 0.0643 0.4913 0.3904 0.4216 0.3307 0.2663 $R^{2}$ 0.9753 0.9822 0.9990 0.9603 0.9754 0.9715 0.9825 0.9880 Laplacian RMSE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0648 0.0492 0.0397 MAE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0463 0.0352 0.0279 $R^{2}$ 1.0000 1.0000 1.0000 1.0000 1.0000 0.9997 0.9998 0.9999 Sigmoid RMSE 1.8283 0.8245 1.8096 1.2541 1.2770 1.0701 1.9611 2.2970 MAE 1.3103 0.5973 1.3112 0.9025 0.9082 0.7856 1.4166 1.7306 $R^{2}$ 0.7473 0.9429 0.7354 0.8782 0.8768 0.9162 0.7237 0.6295 Cosine RMSE 6.4854 7.0431 7.1016 7.1794 7.1538 7.0918 6.9837 6.8637 MAE 5.5633 6.1464 6.1758 6.2249 6.1695 6.1030 5.9616 5.8011 $R^{2}$2.1793 -3.1696 -3.0751 -2.9934 -2.8659 -2.6827 -2.5044 -2.3079
@@
Table VII: Learning performance metrics for the models trained with the hyperparameters listed in Table II for $\mathscr{O}_{Sum}$. The metrics are calculated based on the predictions made by the models on the training and testing sets for $\mathscr{H}_{1}$ with labels determined by $\mathscr{O}_{Sum}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets. All values in this table are plotted in Figure 7, though not all of them are visible within the chosen plot range.
C.3 Results for $\mathscr{O}_{XZ}$ and $\mathscr{H}_{2}$
@@
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear RMSE 0.7436 0.7882 0.8055 0.8159 0.8166 0.8162 0.8148 0.8127 MAE 0.6594 0.7145 0.7385 0.7505 0.7493 0.7454 0.7407 0.7354 $R^{2}$2.1150 -3.8313 -4.6197 -4.9758 -4.8876 -4.6711 -4.4779 -4.2726 Polynomial RMSE 0.2462 0.2088 0.1877 0.1703 0.1717 0.1731 0.1741 0.1683 MAE 0.1804 0.1481 0.1335 0.1248 0.1230 0.1264 0.1278 0.1254 $R^{2}$ 0.6585 0.6609 0.6947 0.7395 0.7397 0.7450 0.7500 0.7739 RBF RMSE 0.0782 0.1308 0.1399 0.1232 0.1202 0.1152 0.1123 0.1065 MAE 0.0457 0.0848 0.0894 0.0865 0.0827 0.0804 0.0740 0.0695 $R^{2}$ 0.9656 0.8670 0.8306 0.8638 0.8725 0.8869 0.8959 0.9095 Laplacian RMSE 0.1355 0.1372 0.1337 0.1179 0.1149 0.1100 0.1110 0.1232 MAE 0.0839 0.0920 0.0858 0.0778 0.0744 0.0709 0.0716 0.0880 $R^{2}$ 0.8965 0.8536 0.8452 0.8751 0.8834 0.8970 0.8982 0.8789 Sigmoid RMSE 0.3297 0.1771 0.2385 0.1724 0.1671 0.1482 0.2161 0.2327 MAE 0.2513 0.1221 0.1712 0.1254 0.1198 0.1063 0.1541 0.1646 $R^{2}$ 0.3878 0.7562 0.5072 0.7333 0.7533 0.8132 0.6145 0.5679 Cosine RMSE 0.7445 0.7888 0.8067 0.8158 0.8165 0.8161 0.8144 0.8122 MAE 0.6603 0.7146 0.7396 0.7502 0.7488 0.7447 0.7400 0.7343 $R^{2}$2.1225 -3.8386 -4.6360 -4.9752 -4.8860 -4.6687 -4.4737 -4.2664 Training Linear RMSE 0.7533 0.7960 0.8159 0.8177 0.8237 0.8204 0.8194 0.8164 MAE 0.6691 0.7256 0.7532 0.7537 0.7588 0.7515 0.7479 0.7414 $R^{2}$2.0465 -3.9309 -5.1218 -5.1428 -5.2954 -4.9662 -4.8482 -4.5928 Polynomial RMSE 0.2321 0.2011 0.1691 0.1580 0.1565 0.1598 0.1589 0.1560 MAE 0.1741 0.1374 0.1204 0.1111 0.1100 0.1115 0.1125 0.1123 $R^{2}$ 0.7107 0.6852 0.7371 0.7708 0.7728 0.7736 0.7800 0.7957 RBF RMSE 0.0035 0.1112 0.0215 0.1018 0.0968 0.0862 0.0702 0.0578 MAE 0.0020 0.0663 0.0124 0.0649 0.0642 0.0556 0.0457 0.0368 $R^{2}$ 0.9999 0.9037 0.9958 0.9047 0.9131 0.9341 0.9571 0.9720 Laplacian RMSE 0.0000 0.0765 0.0087 0.0060 0.0000 0.0000 0.0000 0.0355 MAE 0.0000 0.0485 0.0058 0.0038 0.0000 0.0000 0.0000 0.0259 $R^{2}$ 1.0000 0.9544 0.9993 0.9997 1.0000 1.0000 1.0000 0.9894 Sigmoid RMSE 0.3343 0.1586 0.2264 0.1580 0.1554 0.1277 0.1979 0.2320 MAE 0.2581 0.1065 0.1629 0.1132 0.1097 0.0874 0.1414 0.1670 $R^{2}$ 0.4002 0.8042 0.5287 0.7707 0.7758 0.8554 0.6587 0.5484 Cosine RMSE 0.7530 0.7963 0.8156 0.8183 0.8244 0.8212 0.8202 0.8174 MAE 0.6687 0.7265 0.7530 0.7546 0.7600 0.7529 0.7493 0.7428 $R^{2}$2.0444 -3.9348 -5.1180 -5.1514 -5.3066 -4.9779 -4.8603 -4.6056
Table VIII: Learning performance metrics for the models trained with the hyperparameters listed in Table III for $\mathscr{O}_{XZ}$. The metrics are calculated based on the predictions made by the models on the training and testing sets for $\mathscr{H}_{2}$ with labels determined by $\mathscr{O}_{XZ}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets. All values in this table are plotted in Figure 9, though not all of them are visible within the chosen plot range.
C.4 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{2}$
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear RMSE 6.6084 6.9384 7.1717 7.2117 7.2634 7.2450 7.2467 7.2143 MAE 5.7357 6.2669 6.5641 6.6096 6.6403 6.5921 6.5668 6.5059 $R^{2}$2.4632 -4.3090 -5.1549 -5.2246 -$5$4.7658 -4.5452 -4.3089 Polynomial RMSE 1.4942 1.1867 1.3563 1.2833 1.3342 1.3341 1.7604 1.4036 MAE 1.1004 0.8443 1.0029 0.9740 1.0096 1.0189 1.2957 1.0740 $R^{2}$ 0.8230 0.8447 0.7799 0.8029 0.7957 0.8045 0.6727 0.7990 RBF RMSE 0.6089 0.9284 0.9039 0.9337 0.9487 0.9708 0.9471 0.9099 MAE 0.3170 0.6004 0.5720 0.6676 0.6700 0.6674 0.6390 0.6063 $R^{2}$ 0.9706 0.9049 0.9022 0.8957 0.8967 0.8965 0.9053 0.9156 Laplacian RMSE 1.0664 0.9648 0.9189 0.8730 0.8813 1.0177 1.0370 1.0398 MAE 0.7712 0.6657 0.6309 0.6007 0.5915 0.7304 0.7410 0.7411 $R^{2}$ 0.9098 0.8973 0.8990 0.9088 0.9108 0.8862 0.8865 0.8897 Sigmoid RMSE 2.5193 1.1984 1.9133 1.2459 1.2601 1.2474 1.8645 2.0244 MAE 1.9383 0.8575 1.3799 0.9470 0.9429 0.8848 1.3602 1.4461 $R^{2}$ 0.4967 0.8416 0.5619 0.8142 0.8178 0.8291 0.6329 0.5820 Cosine RMSE 6.6158 6.9427 7.1717 7.2107 7.2616 7.2411 7.2422 7.2080 MAE 5.7417 6.2682 6.5637 6.6062 6.6348 6.5832 6.5569 6.4933 $R^{2}$2.4709 -4.3156 -5.1549 -5.2228 -5.0525 -4.7597 -4.5384 -4.2995 Training Linear RMSE 6.6747 6.9451 7.2232 7.2249 7.2857 7.2512 7.2503 7.2167 MAE 5.7924 6.2699 6.6278 6.6248 6.6670 6.6053 6.5781 6.5165 $R^{2}$2.5404 -4.3985 -5.3317 -5.2921 -5.1667 -4.8959 -4.6703 -4.4283 Polynomial RMSE 1.2201 1.0716 1.3412 1.2795 1.3206 1.3084 1.6460 1.3672 MAE 0.9328 0.7624 0.9493 0.9258 0.9539 0.9551 1.2068 1.0022 $R^{2}$ 0.8817 0.8715 0.7817 0.8027 0.7974 0.8080 0.7078 0.8052 RBF RMSE 0.0252 0.7066 0.1727 0.7663 0.7902 0.6823 0.5885 0.4879 MAE 0.0142 0.4788 0.1020 0.5265 0.5465 0.4763 0.4020 0.3227 $R^{2}$ 0.9999 0.9441 0.9964 0.9292 0.9275 0.9478 0.9626 0.9752 Laplacian RMSE 0.7618 0.4858 0.0639 0.0435 0.0000 0.3533 0.3186 0.2906 MAE 0.5515 0.3406 0.0450 0.0301 0.0000 0.2641 0.2397 0.2191 $R^{2}$ 0.9539 0.9736 0.9995 0.9998 1.0000 0.9860 0.9891 0.9912 Sigmoid RMSE 2.4146 1.0306 1.8288 1.2287 1.2850 1.0282 1.7178 2.0124 MAE 1.8768 0.7487 1.3039 0.8932 0.9275 0.7365 1.2508 1.4609 $R^{2}$ 0.5367 0.8811 0.5941 0.8180 0.8082 0.8815 0.6817 0.5779 Cosine RMSE 6.6723 6.9480 7.2267 7.2301 7.2915 7.2585 7.2572 7.2243 MAE 5.7902 6.2753 6.6324 6.6318 6.6754 6.6159 6.5880 6.5268 $R^{2}$2.5378 -4.4031 -5.3380 -5.3013 -5.1766 -4.9077 -4.6811 -4.4397
@@ C.4 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{2}$
Table IX: Learning performance metrics for the models trained with the hyperparameters listed in Table III for $\mathscr{O}_{Sum}$. The metrics are calculated based on the predictions made by the models on the training and testing sets for $\mathscr{H}_{2}$ with labels determined by $\mathscr{O}_{Sum}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets. All values in this table are plotted in Figure 9, though not all of them are visible within the chosen plot range.
C.5 Results for $\mathscr{O}_{XZ}$ and $\mathscr{H}_{3}$
@@
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear RMSE 0.6230 0.6630 0.6810 0.6884 0.6873 0.6775 0.6611 0.6408 MAE 0.5066 0.5423 0.5542 0.5567 0.5506 0.5349 0.5117 0.4843 $R^{2}$1.5252 -1.9782 -1.9471 -1.8755 -1.7657 -1.6170 -1.4506 -1.2877 Polynomial RMSE 0.1792 0.1395 0.1376 0.1402 0.1392 0.1436 0.1429 0.1480 MAE 0.1323 0.1031 0.1035 0.1066 0.1054 0.1076 0.1059 0.1080 $R^{2}$ 0.7911 0.8681 0.8797 0.8807 0.8866 0.8824 0.8855 0.8780 RBF RMSE 0.0608 0.0502 0.0459 0.0724 0.0655 0.0633 0.0598 0.0564 MAE 0.0379 0.0354 0.0290 0.0511 0.0452 0.0430 0.0384 0.0347 $R^{2}$ 0.9760 0.9829 0.9866 0.9682 0.9749 0.9772 0.9799 0.9823 Laplacian RMSE 0.0843 0.0780 0.0704 0.0679 0.0671 0.0670 0.0675 0.0688 MAE 0.0573 0.0570 0.0498 0.0470 0.0456 0.0445 0.0431 0.0424 $R^{2}$ 0.9537 0.9588 0.9685 0.9720 0.9736 0.9744 0.9745 0.9736 Sigmoid RMSE 0.1965 0.1105 0.1727 0.1417 0.1415 0.1257 0.1085 0.2197 MAE 0.1504 0.0835 0.1336 0.1067 0.1092 0.0975 0.0818 0.1781 $R^{2}$ 0.7489 0.9173 0.8105 0.8782 0.8828 0.9099 0.9340 0.7311 Cosine RMSE 0.6234 0.6635 0.6810 0.6884 0.6873 0.6775 0.6611 0.6407 MAE 0.5068 0.5431 0.5539 0.5561 0.5493 0.5326 0.5087 0.4808 $R^{2}$1.5284 -1.9820 -1.9474 -1.8757 -1.7659 -1.6170 -1.4505 -1.2875 Training Linear RMSE 0.6146 0.6641 0.6847 0.6934 0.6927 0.6825 0.6646 0.6424 MAE 0.5005 0.5452 0.5581 0.5607 0.5547 0.5390 0.5154 0.4871 $R^{2}$1.$5$2.0435 -1.9755 -1.8833 -1.7698 -1.6269 -1.4667 -1.3070 Polynomial RMSE 0.1640 0.1285 0.1270 0.1279 0.1251 0.1293 0.1270 0.1301 MAE 0.1217 0.0954 0.0960 0.0984 0.0952 0.0969 0.0949 0.0966 $R^{2}$ 0.8181 0.8861 0.8976 0.9018 0.9097 0.9057 0.9100 0.9054 RBF RMSE 0.0108 0.0282 0.0118 0.0506 0.0474 0.0377 0.0282 0.0196 MAE 0.0070 0.0190 0.0069 0.0355 0.0322 0.0255 0.0185 0.0123 $R^{2}$ 0.9992 0.9945 0.9991 0.9847 0.9871 0.9920 0.9956 0.9979 Laplacian RMSE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 MAE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 $R^{2}$ 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 Sigmoid RMSE 0.2169 0.0894 0.1814 0.1245 0.1321 0.1071 0.0888 0.2324 MAE 0.1649 0.0693 0.1367 0.0965 0.1015 0.0839 0.0691 0.1849 $R^{2}$ 0.6817 0.9448 0.7911 0.9070 0.8992 0.9353 0.9559 0.6980 Cosine RMSE 0.6143 0.6637 0.6849 0.6936 0.6929 0.6828 0.6649 0.6427 MAE 0.5001 0.5450 0.5582 0.5607 0.5541 0.5375 0.5130 0.4840 $R^{2}$1.5528 -2.0398 -1.9766 -1.8849 -1.7718 -1.6290 -1.4689 -1.3090
@@
Table X: Learning performance metrics for the models trained with the hyperparameters listed in Table IV for $\mathscr{O}_{XZ}$. The metrics are calculated based on the predictions made by the models on the training and testing sets for $\mathscr{H}_{3}$ with labels determined by $\mathscr{O}_{XZ}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets. All values in this table are plotted in Figure 11, though not all of them are visible within the chosen plot range.
C.6 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{3}$

### v2

@@ D.2 Results for $\mathscr{O}_{XZ}$ and $\mathscr{H}_{2}$
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear $R^{2}$2.1150 -3.8313 -4.6197 -4.9758 -4.8876 -4.6711 -4.4779 -4.2726 RMSE 0.7436 0.7882 0.8055 0.8159 0.8166 0.8162 0.8148 0.8127 MAE 0.6594 0.7145 0.7385 0.7505 0.7493 0.7454 0.7407 0.7354 Polynomial $R^{2}$ 0.6585 0.6609 0.6947 0.7395 0.7397 0.7450 0.7500 0.7739 RMSE 0.2462 0.2088 0.1877 0.1703 0.1717 0.1731 0.1741 0.1683 MAE 0.1804 0.1481 0.1335 0.1248 0.1230 0.1264 0.1278 0.1254 RBF $R^{2}$ 0.9656 0.8670 0.8306 0.8638 0.8725 0.8869 0.8959 0.9095 RMSE 0.0782 0.1308 0.1399 0.1232 0.1202 0.1152 0.1123 0.1065 MAE 0.0457 0.0848 0.0894 0.0865 0.0827 0.0804 0.0740 0.0695 Laplacian $R^{2}$ 0.8965 0.8536 0.8452 0.8751 0.8834 0.8970 0.8982 0.8789 RMSE 0.1355 0.1372 0.1337 0.1179 0.1149 0.1100 0.1110 0.1232 MAE 0.0839 0.0920 0.0858 0.0778 0.0744 0.0709 0.0716 0.0880 Sigmoid $R^{2}$ 0.3878 0.7562 0.5072 0.7333 0.7533 0.8132 0.6145 0.5679 RMSE 0.3297 0.1771 0.2385 0.1724 0.1671 0.1482 0.2161 0.2327 MAE 0.2513 0.1221 0.1712 0.1254 0.1198 0.1063 0.1541 0.1646 Cosine $R^{2}$2.1225 -3.8386 -4.6360 -4.9752 -4.8860 -4.6687 -4.4737 -4.2664 RMSE 0.7445 0.7888 0.8067 0.8158 0.8165 0.8161 0.8144 0.8122 MAE 0.6603 0.7146 0.7396 0.7502 0.7488 0.7447 0.7400 0.7343 Training Linear $R^{2}$2.0465 -3.9309 -5.1218 -5.1428 -5.2954 -4.9662 -4.8482 -4.5928 RMSE 0.7533 0.7960 0.8159 0.8177 0.8237 0.8204 0.8194 0.8164 MAE 0.6691 0.7256 0.7532 0.7537 0.7588 0.7515 0.7479 0.7414 Polynomial $R^{2}$ 0.7107 0.6852 0.7371 0.7708 0.7728 0.7736 0.7800 0.7957 RMSE 0.2321 0.2011 0.1691 0.1580 0.1565 0.1598 0.1589 0.1560 MAE 0.1741 0.1374 0.1204 0.1111 0.1100 0.1115 0.1125 0.1123 RBF $R^{2}$ 0.9999 0.9037 0.9958 0.9047 0.9131 0.9341 0.9571 0.9720 RMSE 0.0035 0.1112 0.0215 0.1018 0.0968 0.0862 0.0702 0.0578 MAE 0.0020 0.0663 0.0124 0.0649 0.0642 0.0556 0.0457 0.0368 Laplacian $R^{2}$ 1.0000 0.9544 0.9993 0.9997 1.0000 1.0000 1.0000 0.9894 RMSE 0.0000 0.0765 0.0087 0.0060 0.0000 0.0000 0.0000 0.0355 MAE 0.0000 0.0485 0.0058 0.0038 0.0000 0.0000 0.0000 0.0259 Sigmoid $R^{2}$ 0.4002 0.8042 0.5287 0.7707 0.7758 0.8554 0.6587 0.5484 RMSE 0.3343 0.1586 0.2264 0.1580 0.1554 0.1277 0.1979 0.2320 MAE 0.2581 0.1065 0.1629 0.1132 0.1097 0.0874 0.1414 0.1670 Cosine $R^{2}$2.0444 -3.9348 -5.1180 -5.1514 -5.3066 -4.9779 -4.8603 -4.6056 RMSE 0.7530 0.7963 0.8156 0.8183 0.8244 0.8212 0.8202 0.8174 MAE 0.6687 0.7265 0.7530 0.7546 0.7600 0.7529 0.7493 0.7428
@@
Table D.2: ($\mathscr{O}_{XZ}$ with $\mathscr{H}_{2}$) Learning performance metrics for the models trained on all 1000 training datapoints using the hyperparameters listed in Table C.3. The metrics are calculated based on the predictions made by the models on the training and testing sets with labels determined by $\mathscr{O}_{XZ}$ with $\mathscr{H}_{2}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets.
D.3 Results for $\mathscr{O}_{XZ}$ and $\mathscr{H}_{3}$
@@
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear $R^{2}$1.5252 -1.9782 -1.9471 -1.8755 -1.7657 -1.6170 -1.4506 -1.2877 RMSE 0.6230 0.6630 0.6810 0.6884 0.6873 0.6775 0.6611 0.6408 MAE 0.5066 0.5423 0.5542 0.5567 0.5506 0.5349 0.5117 0.4843 Polynomial $R^{2}$ 0.7911 0.8681 0.8797 0.8807 0.8866 0.8824 0.8855 0.8780 RMSE 0.1792 0.1395 0.1376 0.1402 0.1392 0.1436 0.1429 0.1480 MAE 0.1323 0.1031 0.1035 0.1066 0.1054 0.1076 0.1059 0.1080 RBF $R^{2}$ 0.9760 0.9829 0.9866 0.9682 0.9749 0.9772 0.9799 0.9823 RMSE 0.0608 0.0502 0.0459 0.0724 0.0655 0.0633 0.0598 0.0564 MAE 0.0379 0.0354 0.0290 0.0511 0.0452 0.0430 0.0384 0.0347 Laplacian $R^{2}$ 0.9537 0.9588 0.9685 0.9720 0.9736 0.9744 0.9745 0.9736 RMSE 0.0843 0.0780 0.0704 0.0679 0.0671 0.0670 0.0675 0.0688 MAE 0.0573 0.0570 0.0498 0.0470 0.0456 0.0445 0.0431 0.0424 Sigmoid $R^{2}$ 0.7489 0.9173 0.8105 0.8782 0.8828 0.9099 0.9340 0.7311 RMSE 0.1965 0.1105 0.1727 0.1417 0.1415 0.1257 0.1085 0.2197 MAE 0.1504 0.0835 0.1336 0.1067 0.1092 0.0975 0.0818 0.1781 Cosine $R^{2}$1.5284 -1.9820 -1.9474 -1.8757 -1.7659 -1.6170 -1.4505 -1.2875 RMSE 0.6234 0.6635 0.6810 0.6884 0.6873 0.6775 0.6611 0.6407 MAE 0.5068 0.5431 0.5539 0.5561 0.5493 0.5326 0.5087 0.4808 Training Linear $R^{2}$1.$5$2.0435 -1.9755 -1.8833 -1.7698 -1.6269 -1.4667 -1.3070 RMSE 0.6146 0.6641 0.6847 0.6934 0.6927 0.6825 0.6646 0.6424 MAE 0.5005 0.5452 0.5581 0.5607 0.5547 0.5390 0.5154 0.4871 Polynomial $R^{2}$ 0.8181 0.8861 0.8976 0.9018 0.9097 0.9057 0.9100 0.9054 RMSE 0.1640 0.1285 0.1270 0.1279 0.1251 0.1293 0.1270 0.1301 MAE 0.1217 0.0954 0.0960 0.0984 0.0952 0.0969 0.0949 0.0966 RBF $R^{2}$ 0.9992 0.9945 0.9991 0.9847 0.9871 0.9920 0.9956 0.9979 RMSE 0.0108 0.0282 0.0118 0.0506 0.0474 0.0377 0.0282 0.0196 MAE 0.0070 0.0190 0.0069 0.0355 0.0322 0.0255 0.0185 0.0123 Laplacian $R^{2}$ 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 RMSE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 MAE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 Sigmoid $R^{2}$ 0.6817 0.9448 0.7911 0.9070 0.8992 0.9353 0.9559 0.6980 RMSE 0.2169 0.0894 0.1814 0.1245 0.1321 0.1071 0.0888 0.2324 MAE 0.1649 0.0693 0.1367 0.0965 0.1015 0.0839 0.0691 0.1849 Cosine $R^{2}$1.5528 -2.0398 -1.9766 -1.8849 -1.7718 -1.6290 -1.4689 -1.3090 RMSE 0.6143 0.6637 0.6849 0.6936 0.6929 0.6828 0.6649 0.6427 MAE 0.5001 0.5450 0.5582 0.5607 0.5541 0.5375 0.5130 0.4840
@@
Table D.3: ($\mathscr{O}_{XZ}$ with $\mathscr{H}_{3}$) Learning performance metrics for the models trained on all 1000 training datapoints using the hyperparameters listed in Table C.4. The metrics are calculated based on the predictions made by the models on the training and testing sets with labels determined by $\mathscr{O}_{XZ}$ with $\mathscr{H}_{3}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets.
D.4 Results for $\mathscr{O}_{XZ}$ and $\mathscr{H}_{4}$
@@
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear $R^{2}$5.4877 -9.7666 -12.2134 -13.2260 -13.3083 -12.7522 -11.8802 -10.6947 RMSE 0.8851 0.9240 0.9337 0.9401 0.9424 0.9408 0.9360 0.9293 MAE 0.8375 0.8844 0.8976 0.9061 0.9084 0.9053 0.8982 0.8878 Polynomial $R^{2}$ 0.7034 0.6119 0.6135 0.6729 0.6911 0.7070 0.7900 0.7817 RMSE 0.1892 0.1754 0.1597 0.1425 0.1385 0.1373 0.1195 0.1270 MAE 0.1060 0.0914 0.0822 0.0782 0.0745 0.0814 0.0690 0.0742 RBF $R^{2}$ 0.8360 0.7167 0.7751 0.8281 0.7555 0.8177 0.8674 0.8931 RMSE 0.1407 0.1499 0.1218 0.1033 0.1232 0.1083 0.0950 0.0888 MAE 0.0724 0.0731 0.0522 0.0498 0.0650 0.0577 0.0507 0.0468 Laplacian $R^{2}$ 0.9214 0.8612 0.7511 0.7732 0.7835 0.8140 0.8521 0.8817 RMSE 0.0974 0.1049 0.1282 0.1187 0.1159 0.1094 0.1003 0.0934 MAE 0.0514 0.0521 0.0715 0.0686 0.0658 0.0623 0.0575 0.0536 Sigmoid $R^{2}$ 0.6985 0.6030 0.6269 0.4990 0.5695 -3.7352 0.3487 -11636.6066 RMSE 0.1908 0.1774 0.1569 0.1764 0.1635 0.5520 0.2105 29.3142 MAE 0.0981 0.0940 0.0815 0.1014 0.1055 0.3001 0.1254 17.0978 Cosine $R^{2}$5.4872 -9.7161 -12.2097 -13.1944 -13.2552 -12.6843 -11.8028 -10.6139 RMSE 0.8851 0.9218 0.9336 0.9391 0.9407 0.9384 0.9332 0.9261 MAE 0.8376 0.8821 0.8973 0.9050 0.9068 0.9033 0.8959 0.8851 Training Linear $R^{2}$6.5416 -11.1825 -12.0591 -12.1214 -11.7599 -11.1197 -10.3358 -9.4707 RMSE 0.8912 0.9224 0.9317 0.9352 0.9363 0.9345 0.9300 0.9233 MAE 0.8470 0.8860 0.8965 0.9001 0.9002 0.8968 0.8903 0.8810 Polynomial $R^{2}$ 0.7641 0.8003 0.8095 0.8358 0.8196 0.8226 0.8916 0.8879 RMSE 0.1576 0.1181 0.1125 0.1046 0.1113 0.1131 0.0909 0.0955 MAE 0.0884 0.0656 0.0616 0.0616 0.0641 0.0704 0.0514 0.0555 RBF $R^{2}$ 0.9374 0.9012 0.9703 0.9927 0.8783 0.9120 0.9443 0.9666 RMSE 0.0812 0.0831 0.0444 0.0221 0.0914 0.0796 0.0652 0.0521 MAE 0.0434 0.0440 0.0226 0.0103 0.0510 0.0429 0.0349 0.0275 Laplacian $R^{2}$ 1.0000 1.0000 0.9991 0.9996 0.9998 0.9999 1.0000 1.0000 RMSE 0.0000 0.0000 0.0076 0.0048 0.0034 0.0026 0.0019 0.0002 MAE 0.0000 0.0000 0.0051 0.0032 0.0022 0.0016 0.0012 0.0001 Sigmoid $R^{2}$ 0.8788 0.7988 0.8540 0.5826 0.6472 -4.3227 0.4554 -13333.5415 RMSE 0.1130 0.1185 0.0985 0.1668 0.1557 0.6193 0.2038 32.9478 MAE 0.0658 0.0673 0.0550 0.0976 0.0989 0.3556 0.1172 20.1928 Cosine $R^{2}$6.5589 -11.2173 -12.0804 -12.1504 -$1$11.1499 -10.3634 -9.4954 RMSE 0.8922 0.9237 0.9325 0.9363 0.9374 0.9357 0.9311 0.9243 MAE 0.8483 0.8874 0.8969 0.9003 0.9003 0.8967 0.8896 0.8797
@@
Table D.4: ($\mathscr{O}_{XZ}$ with $\mathscr{H}_{4}$) Learning performance metrics for the models trained on all 1000 training datapoints using the hyperparameters listed in Table C.5. The metrics are calculated based on the predictions made by the models on the training and testing sets with labels determined by $\mathscr{O}_{XZ}$ with $\mathscr{H}_{4}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets.
D.5 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{1}$